# Schrödinger cats escape confinement

### v2

@@ Acknowledgments
We are grateful to M.A. Werner for sharing his extensive knowledge of MPS methods and helping us develop our simulations. We also thank M. Kormos and M. Lencsés for discussions, and the latter especially for suggesting the title of our work. We are grateful to A. Bastianello for useful comments on the initial version of the manuscript. This work was supported by the National Research, Development and Innovation Office (NKFIH) through the OTKA Grants SNN 139581 and K 138606. AK was also partially supported by the Doctoral Excellence Fellowship Programme (DCEP), funded by the National Research Development and Innovation Fund of the Ministry of Culture and Innovation and the Budapest University of Technology and Economics, under a grant agreement with the National Research, Development and Innovation Office. GT was partially supported by the Quantum Information National Laboratory of Hungary (Grant No. 2022-2.1.1-NL-2022-00004).

### v3

@@ Acknowledgments
We are grateful to M.A. Werner for sharing his extensive knowledge of MPS methods and helping us develop our simulations. We also thank M. Kormos and M. Lencsés for discussions, and are grateful to A. Bastianello for useful comments on the initial version of the manuscript. This work was supported by the National Research, Development and Innovation Office (NKFIH) through the OTKA Grants SNN 139581 and K 138606. AK was also partially supported by the Doctoral Excellence Fellowship Programme (DCEP), funded by the National Research Development and Innovation Fund of the Ministry of Culture and Innovation and the Budapest University of Technology and Economics, under a grant agreement with the National Research, Development and Innovation Office. GT was partially supported by the Quantum Information National Laboratory of Hungary (Grant No. 2022-2.1.1-NL-2022-00004).
@@ Appendix B Numerical results for meson and bubble wave functions
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption